Reeve Corporation has total fixed costs of $275,000.Product X has a selling price of $74.00 per unit and a variable cost of $54.00 per unit.

Required: Calculate the breakeven point in units.
Answer #1

Contribution margin=Sales-Variable cost  

=74-54=$20 per unit

Breakeven=Fixed expenses/Contribution margin

=275,000/20

=13750 units